OpenStudy (anonymous):

I am having a hard time with these finding the equation of the line containing the given pairs of points. Heres another one with fractions which I am horrible at: (1/6, -1/3) and (5/6, 1) Thanks again!

OpenStudy (anonymous):

plug in the points to the equation to obtain the slope of the line which both points are on m = (y2-y1)/(x2-x1) m= (1-(-1/3)/(5/6-1/6) simplifying you get m=2 then use the equation of a line y=mx+b plug in one of the two points and your m to find b 1=2(5/6) + b thus b= -2/3 now for the general equation of a line, you plug in m and b y=2x-(2/3)

OpenStudy (anonymous):

Thank you! You are very good at those!